# Optimal Convergence Rate for Mirror Descent Methods with special Time-Varying Step Sizes Rules

### v3

@@ 4. Mirror descent method for constrained composite problem
    $\operatorname{dom}(h)\subseteq\operatorname{int}(\operatorname{dom}(f))$;
@@
Assume that $b\in\operatorname{dom}(\partial(\psi))$, and let $a$ be defined by
@@
As a special case of Theorem 4.3, with a special value of the parameter $m$, we can deduce the well-known sub-optimal convergence rate of the mirror-C descent method (Algorithm 2), namely $O\left(\frac{\log(N)}{N}\right)$, with the time-varying step sizes given in (3.7) (see [4, 12, 20, 25]).
@@
Suppose that Assumption 4.1 holds and $h$ is a non-negative function. Then for problem (4.1), by Algorithm 2, with $m=-1$, and the time-varying step sizes given in (3.7), it satisfies the following sub-optimal convergence rate
(4.12) $F(\tilde{x})-F(x^{*})$ $\leq\frac{M_{f}}{\sqrt{\sigma}}\left(\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+V_{\psi}(x^{*},x^{1})+1+\log(N)\right)\frac{1}{\sqrt{N}}$
@@
When $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}$, then by substitution in (4.13), with $\tilde{x}=\frac{1}{\sum_{k=1}^{N}\frac{1}{\sqrt{k}}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}x^{k}$, we find
$F(\tilde{x})-F(x^{*})$ $\leq\frac{M_{f}}{\sqrt{2\sigma}}\frac{\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+V_{\psi}(x^{*},x^{1})+\sum_{k=1}^{N}\frac{1}{k}}{\sum_{k=1}^{N}\frac{1}{\sqrt{k}}}$
$\leq\frac{M_{f}}{\sqrt{2\sigma}}\frac{\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+V_{\psi}(x^{*},x^{1})+1+\log(N)}{2\sqrt{N+1}-2}$
$\leq\frac{M_{f}}{\sqrt{\sigma}}\frac{\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+V_{\psi}(x^{*},x^{1})+1+\log(N)}{\sqrt{N}}.$
@@
Case 2 (adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}$, then by substitution in (4.13), we find
$F(\tilde{x})-F(x^{*})$ $\leq\frac{\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+V_{\psi}(x^{*},x^{1})+\sum_{k=1}^{N}\frac{1}{k}}{\sum_{k=1}^{N}\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}}}$
$\leq\frac{M_{f}}{\sqrt{\sigma}}\frac{\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+V_{\psi}(x^{*},x^{1})+1+\log(N)}{\sqrt{N}},$
@@
Suppose that Assumption 4.1 hold and $h$ is a non-negative function, with $h(x^{1})=0$. Then for problem (4.1), by Algorithm 2, with $m=0$, and the time-varying step sizes given in (3.7), it satisfies the following optimal convergence rate
$F(\overline{x})-F(x^{*})\leq\left(\frac{M_{f}V_{\psi}(x^{*},x^{1})+2M_{f}}{\sqrt{2\sigma}}\right)\frac{1}{\sqrt{N}}=O\left(\frac{1}{\sqrt{N}}\right),$
@@
Case 1 (non-adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}$, then by substitution in (4.14) we find
$F(\overline{x})-F(x^{*})$ $\leq\frac{1}{N}\left(\frac{M_{f}\sqrt{N}V_{\psi}(x^{*},x^{1})}{\sqrt{2\sigma}}+\frac{M_{f}}{\sqrt{2\sigma}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}\right)$
$\leq\frac{1}{N}\frac{M_{f}}{\sqrt{2\sigma}}\left(\sqrt{N}V_{\psi}(x^{*},x^{1})+2\sqrt{N}\right)$
$=\frac{M_{f}(2+V_{\psi}(x^{*},x^{1}))}{\sqrt{2\sigma}}\frac{1}{\sqrt{N}}.$
Case 2 (adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}$, then by substitution in (4.14) we find
$F(\overline{x})-F(x^{*})$ $\leq\frac{1}{N}\left(\frac{M_{f}\sqrt{N}V_{\psi}(x^{*},x^{1})}{\sqrt{2\sigma}}+\frac{M_{f}}{\sqrt{2\sigma}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}\right)$
$\leq\frac{M_{f}(2+V_{\psi}(x^{*},x^{1}))}{\sqrt{2\sigma}}\frac{1}{\sqrt{N}}.$
@@
By setting any $m>-1$ in (4.3) with the assumption that $h(x^{1})=0$, the convergence rate of Algorithm 2 with the time-varying step sizes given in (3.7), will be optimal without the presence of a $\log(N)$ factor, namely $O\left(\frac{1}{\sqrt{N}}\right)$. In comparison with the sub-optimal convergence rate (4.12), when $m>0$, the weighting scheme $\frac{1}{\sum_{k=1}^{N}\gamma_{k}^{-m}}\sum_{k=1}^{N}\gamma_{k}^{-m}x^{k}$ assigns smaller weights to the initial points and larger weights to the most recent points that generated by Algorithm 2.
@@ 5. Mirror descent method for problems with functional constraints
where $f$ and $g_{i}$ (for all $i=1,\ldots,p$) are convex and non-smooth functions given on a closed convex set $Q$.
It is clear that instead of a set of functionals $\{g_{i}(\cdot)\}_{i=1}^{p}$ we can see one functional constraint $g:Q\rightarrow\mathbb{R}$, such that $g(x)=\max_{1\leq i\leq p}\{g_{i}(x)\}$. Therefore, by this setting, the problem (5.1) will be equivalent to the following constrained minimization problem
@@
The needed point (output of Algorithm 3) will be selected among the points $x^{k}$ for which $g(x^{k})\leq\varepsilon$. Therefore, we will call step $k$ productive if $g(x^{k})\leq\varepsilon$. If the reverse inequality $g(x^{k})\geq\varepsilon$ holds then step $k$ will be called non-productive. Let $I$ and $J$ denote the set of productive and non-productive steps, respectively. $|I|$ and $|J|$ denote the number of productive and non-productive steps, respectively. Let us also set $\gamma_{k}:=\gamma_{k}^{f}$ if $k\in I$ and $\gamma_{k}:=\gamma_{k}^{g}$ if $k\in J$.
@@
where $\hat{x}=\frac{1}{\sum_{k\in I}\left(\gamma_{k}^{f}\right)^{-m}}\sum_{k\in I}\left(\gamma_{k}^{f}\right)^{-m}x^{k}$ and $x^{*}$ is an optimal solution of (5.2).
@@
(5.5) $\frac{g(x^{k})-g(x^{*})}{\left(\gamma_{k}^{g}\right)^{m}}\leq\frac{V_{\psi}(x^{*},x^{k})-V_{\psi}(x^{*},x^{k+1})}{\left(\gamma_{k}^{g}\right)^{m+1}}+\frac{\|\nabla g(x^{k})\|_{*}^{2}}{2\sigma\left(\gamma_{k}^{g}\right)^{m-1}},$
@@
Also, in a similar way as in the proof of Theorem 3.1, and since for any $k\in J$, we have
(5.6) $g(x^{k})-g(x^{*})\geq g(x^{k})>\varepsilon>0.$
Therefore we get
@@
By dividing both sides of the previous inequality by $\sum_{k\in I}\left(\gamma_{k}^{f}\right)^{-m}$, we get the desired inequality (5.3).
Finally, let us show that $|I|\neq 0$. For this let us assume that $|I|=0$, therefore $|J|=N$, i.e., all steps are non-productive. From (5.6), we get
$\sum_{k=1}^{N}\frac{g(x^{k})-g(x^{*})}{\gamma_{k}^{m}}>\sum_{k=1}^{N}\frac{\varepsilon}{\gamma_{k}^{m}}=\frac{\varepsilon M^{m}}{\left(\sqrt{2\sigma}\right)^{m}}\sum_{k=1}^{N}\left(\sqrt{k}\right)^{m},$
and for all $k\in J=\{1,\ldots,N\}$, with the assumption that $V_{\psi}(x^{*},x^{1})\leq\theta$ for some $\theta>0$, we get
@@
So, we have a contradiction. It means that $|I|\neq 0$. ∎
@@
From inequality (5.7) (which is equivalent to the inequality (5.3)), with $\gamma_{q}=\gamma_{q}^{f}$ if $q\in I$ and $\gamma_{q}=\gamma_{q}^{g}$ if $q\in J$, with
@@
$\quad=\varepsilon\sum_{i\in I}\left(\gamma_{i}^{f}\right)^{-m}-\Bigg{(}\varepsilon\sum_{i=1}^{k}\left(\gamma_{i}\right)^{-m}-\frac{V_{\psi}(x^{*},x^{1})}{\gamma_{k}^{m+1}}-\frac{1}{2\sigma}\sum_{i\in I}\frac{\|\nabla f(x^{i})\|_{*}^{2}}{\left(\gamma_{i}^{f}\right)^{m-1}}$
@@
(5.10) $\varepsilon\sum_{i=1}^{k}\left(\gamma_{i}\right)^{-m}\geq\frac{V_{\psi}(x^{*},x^{1})}{\gamma_{k}^{m+1}}+\frac{1}{2\sigma}\sum_{i\in I}\frac{\|\nabla f(x^{i})\|_{*}^{2}}{\left(\gamma_{i}^{f}\right)^{m-1}}+\frac{1}{2\sigma}\sum_{j\in J}\frac{\|\nabla g(x^{j})\|_{*}^{2}}{\left(\gamma_{j}^{g}\right)^{m-1}},$
as a stopping criterion of Algorithm 3 and as a result we conclude
@@
Thus after the stop criterion (5.10) is met we find that $\hat{x}$ given in (5.9) represents an $\varepsilon$-solution to problem (5.2).
@@
(5.12) $N=\left\lceil\frac{M^{2}(1+\theta)^{2}}{2\sigma\varepsilon^{2}}\right\rceil=O\left(\frac{1}{\varepsilon^{2}}\right),$
@@
[…]
@@
For solving problem (5.1) (or its equivalence (5.2)), when we have a very big number of functional constraints, we can modify Algorithm 3 as in [32]. When we have a non-productive step $k$, i.e., $g(x^{k})>\varepsilon$, then instead of calculating a (sub)gradient of the functional constraint with max-type $g(x)=\max_{1\leq i\leq p}\{g_{i}(x^{k})\}$, we calculate a (sub)gradient of one functional $g_{i}$, for which we have $g_{i}(x^{k})>\varepsilon$. This idea of the proposed modification of Algorithms 3 allows saving the running time of the algorithm due to consideration of not all functional constraints on non-productive steps.

### v5

@@ 4. Mirror descent method for constrained composite problem
    $\operatorname{dom}(h)\subseteq\operatorname{int}(\operatorname{dom}(f))$, where $\operatorname{int}(A)$ denotes the interior of the set $A$;
@@
Assume that $b\in\operatorname{dom}(\partial\psi)$, and let $a$ be defined by
@@
As a special case of Theorem 4.3, with a special value of the parameter $m$, we can deduce the well-known sub-optimal convergence rate of the mirror-C descent method (Algorithm 2), namely $O\left(\frac{\log(N)}{\sqrt{N}}\right)$, with the time-varying step sizes given in (3.7) (see [4, 12, 19, 24]).
@@
Suppose that Assumption 4.1 holds and $h$ is a non-negative function. Then for problem (4.1), by Algorithm 2, with $m=-1$, $V_{\psi}(x^{*},x^{1})\leq\theta$, for some $\theta>0$, and with the time-varying step sizes given in (3.7), it satisfies the following sub-optimal convergence rate
(4.12) $F(\tilde{x})-F(x^{*})$ $\leq\frac{M_{f}}{\sqrt{\sigma}}\left(\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+\theta+1+\log(N)\right)\frac{1}{\sqrt{N}}$
@@
When $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f},V_{\psi}(x^{*},x^{1})\leq\theta$, then by substitution in (4.13), with $\tilde{x}=\frac{1}{\sum_{k=1}^{N}\frac{1}{\sqrt{k}}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}x^{k}$, we find
$F(\tilde{x})-F(x^{*})$ $\leq\frac{M_{f}}{\sqrt{2\sigma}}\frac{\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+\theta+\sum_{k=1}^{N}\frac{1}{k}}{\sum_{k=1}^{N}\frac{1}{\sqrt{k}}}$
$\leq\frac{M_{f}}{\sqrt{2\sigma}}\frac{\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+1+\theta+\log(N)}{2\sqrt{N+1}-2}$
$\leq\frac{M_{f}}{\sqrt{\sigma}}\frac{\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+1+\theta+\log(N)}{\sqrt{N}}.$
@@
Case 2 (adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f},V_{\psi}(x^{*},x^{1})\leq\theta$, then by substitution in (4.13), we find
$F(\tilde{x})-F(x^{*})$ $\leq\frac{\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+\theta+\sum_{k=1}^{N}\frac{1}{k}}{\sum_{k=1}^{N}\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}}}$
$\leq\frac{M_{f}}{\sqrt{\sigma}}\frac{\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+1+\theta+\log(N)}{\sqrt{N}},$
@@
Suppose that Assumption 4.1 hold and $h$ is a non-negative function, with $h(x^{1})=0$. Then for problem (4.1), by Algorithm 2, with $m=0$, $V_{\psi}(x^{*},x^{1})\leq\theta$, for some $\theta>0$, and with the time-varying step sizes given in (3.7), it satisfies the following optimal convergence rate
$F(\overline{x})-F(x^{*})\leq\frac{(2+\theta)M_{f}}{\sqrt{2\sigma}}\cdot\frac{1}{\sqrt{N}}=O\left(\frac{1}{\sqrt{N}}\right),$
@@
Case 1 (non-adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f},V_{\psi}(x^{*},x^{1})\leq\theta$, then by substitution in (4.14) we find
$F(\overline{x})-F(x^{*})$ $\leq\frac{1}{N}\left(\frac{\theta M_{f}\sqrt{N}}{\sqrt{2\sigma}}+\frac{M_{f}}{\sqrt{2\sigma}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}\right)\leq\frac{1}{N}\frac{M_{f}}{\sqrt{2\sigma}}\left(\theta\sqrt{N}+2\sqrt{N}\right)$
$=\frac{M_{f}(2+\theta)}{\sqrt{2\sigma}}\cdot\frac{1}{\sqrt{N}}.$
Case 2 (adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f},V_{\psi}(x^{*},x^{1})\leq\theta$, then by substitution in (4.14) we find
$F(\overline{x})-F(x^{*})\leq\frac{1}{N}\left(\frac{\theta M_{f}\sqrt{N}}{\sqrt{2\sigma}}+\frac{M_{f}}{\sqrt{2\sigma}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}\right)\leq\frac{M_{f}(2+\theta)}{\sqrt{2\sigma}}\cdot\frac{1}{\sqrt{N}}.$
@@
In a similar way as in Corollary 3.4, for any $m\geq 1$ in (4.3) with the assumption that $h(x^{1})=0$ we get the same result (3.12), and the convergence rate of Algorithm 2 with the time-varying step sizes given in (3.7), will be optimal. In comparison with the sub-optimal convergence rate (4.12), when $m\geq 1$, the weighting scheme $\frac{1}{\sum_{k=1}^{N}\gamma_{k}^{-m}}\sum_{k=1}^{N}\gamma_{k}^{-m}x^{k}$ assigns smaller weights to the initial points and larger weights to the most recent points that generated by Algorithm 2.
@@ 5. Mirror descent method for problems with functional constraints
where $f$ and $g_{i}$ (for all $i=1,\ldots,p$) are convex and non-smooth functions given on a compact convex set $Q$.
It is clear that instead of a set of functionals $\{g_{i}(\cdot)\}_{i=1}^{p}$ we can see one functional constraint $g:Q\longrightarrow\mathbb{R}$, such that $g(x)=\max_{1\leq i\leq p}\{g_{i}(x)\}$. Therefore, by this setting, the problem (5.1) will be equivalent to the following constrained minimization problem
@@
The output point (output of Algorithm 3) will be selected among the points $x^{k}$ for which $g(x^{k})\leq\varepsilon$. Therefore, we will call step $k$ productive if $g(x^{k})\leq\varepsilon$. If the reverse inequality $g(x^{k})>\varepsilon$ holds then step $k$ will be called non-productive. Let $I$ and $J$ denote the set of productive and non-productive steps, respectively. $|I|$ and $|J|$ denote the number of productive and non-productive steps, respectively. Let us also set $\gamma_{k}:=\gamma_{k}^{f}$ if $k\in I$, $\gamma_{k}:=\gamma_{k}^{g}$ if $k\in J$, and $M=\max\{M_{f},M_{g}\}$.
@@
where $\hat{x}=\frac{1}{\sum_{k\in I}\left(\gamma_{k}^{f}\right)^{-m}}\sum_{k\in I}\left(\gamma_{k}^{f}\right)^{-m}x^{k}$, and $x^{*}$ is an optimal solution of (5.2).
@@
(5.5) $\frac{g(x^{k})-g(x^{*})}{\left(\gamma_{k}^{g}\right)^{m}}\leq\frac{V_{\psi}(x^{*},x^{k})-V_{\psi}(x^{*},x^{k+1})}{\left(\gamma_{k}^{g}\right)^{m+1}}+\frac{\|\nabla g(x^{k})\|_{*}^{2}}{2\sigma\left(\gamma_{k}^{g}\right)^{m-1}}.$
@@
Also, in a similar way as in the proof of Theorem 3.1, and since for any $k\in J$, it holds
(5.6) $g(x^{k})-g(x^{*})\geq g(x^{k})>\varepsilon>0,$
we get
@@
By dividing both sides of (5.7) by $\sum_{k\in I}\left(\gamma_{k}^{f}\right)^{-m}$, we get the desired inequality (5.3).
Now, let us show that $|I|\neq 0$. For this, let us assume that $|I|=0$, therefore $|J|=N$, i.e., all steps are non-productive. From (5.6), we get
(5.8) $\sum_{k=1}^{N}\frac{g(x^{k})-g(x^{*})}{\gamma_{k}^{m}}>\sum_{k=1}^{N}\frac{\varepsilon}{\gamma_{k}^{m}}=\frac{\varepsilon M^{m}}{\left(\sqrt{2\sigma}\right)^{m}}\sum_{k=1}^{N}\left(\sqrt{k}\right)^{m},$
and for all $k\in J=\{1,\ldots,N\}$ with the assumption that $V_{\psi}(x^{*},x^{1})\leq\theta$ for some $\theta>0$, we get
@@
Therefore, we get
$\sum_{k=1}^{N}\frac{g(x^{k})-g(x^{*})}{\gamma_{k}^{m}}<\frac{\varepsilon M^{m}}{\left(\sqrt{2\sigma}\right)^{m}}\sum_{k=1}^{N}\left(\sqrt{k}\right)^{m}.$
So, we have a contradiction with (5.8). This means that $|I|\neq 0$. ∎
@@
Let $x^{1}\in Q$ be an initial point of Algorithm 3, such that $V_{\psi}(x^{*},x^{1})\leq\theta$ for some $\theta>0$. From inequality (5.7) (which is equivalent to the inequality (5.3)), with $\gamma_{q}=\gamma_{q}^{f}$ if $q\in I$ and $\gamma_{q}=\gamma_{q}^{g}$ if $q\in J$, with
@@
$\quad\qquad\leq\varepsilon\sum_{i\in I}\left(\gamma_{i}^{f}\right)^{-m}-\Bigg{(}\varepsilon\sum_{i=1}^{k}\left(\gamma_{i}\right)^{-m}-\frac{\theta}{\gamma_{k}^{m+1}}-\frac{1}{2\sigma}\sum_{i\in I}\frac{\|\nabla f(x^{i})\|_{*}^{2}}{\left(\gamma_{i}^{f}\right)^{m-1}}$
@@
(5.11) $\varepsilon\sum_{i=1}^{k}\left(\gamma_{i}\right)^{-m}\geq\frac{\theta}{\gamma_{k}^{m+1}}+\frac{1}{2\sigma}\sum_{i\in I}\frac{\|\nabla f(x^{i})\|_{*}^{2}}{\left(\gamma_{i}^{f}\right)^{m-1}}+\frac{1}{2\sigma}\sum_{j\in J}\frac{\|\nabla g(x^{j})\|_{*}^{2}}{\left(\gamma_{j}^{g}\right)^{m-1}},$
as a stopping criterion of Algorithm 3. As a result, we conclude
@@
Thus after the stopping criterion (5.11) is met we find that $\hat{x}$, which is given in (5.10), represents an $\varepsilon$-solution to problem (5.2).
@@
(5.13) $N=\left\lceil\frac{M^{2}(1+\theta)^{2}}{2\sigma\varepsilon^{2}}\right\rceil=O\left(\frac{1}{\varepsilon^{2}}\right)$
@@
$f(\overline{x})-f(x^{*})<\varepsilon,\quad\text{and}\quad g(\overline{x})\leq\varepsilon,$
@@
For solving problem (5.1) (or its equivalence (5.2)), when we have a very big number of functional constraints, we can modify Algorithm 3 in a similar way as in [31]. When we have a non-productive step $k$, i.e., it holds that $g(x^{k})>\varepsilon$, then instead of calculating a (sub)gradient of the functional constraint with max-type $g(x)=\max_{1\leq i\leq p}\{g_{i}(x)\}$, we calculate a (sub)gradient of one functional $g_{i}$, for which we have $g_{i}(x^{k})>\varepsilon$. This idea of the modification of Algorithms 3 allows saving the running time of the algorithm due to consideration of not all functional constraints (when their number is big) on non-productive steps.